OpenStudy (anonymous):

Given parametric equations, how can we determine if two lines are perpendicular?

OpenStudy (anonymous):

dot product of their slope is 0

OpenStudy (anonymous):

Thanks you! And by slopes do you mean r'(t)?

OpenStudy (anonymous):

x= 2t + 3 y= 3t+2 z= 2 +3 slope of that line is <2,3,2>

OpenStudy (anonymous):

z=2t+3 ** should be